OpenStudy (anonymous):

Find two negative numbers, x and y, with a product of 70 and a sum which is a maximum. Find two negative numbers, x and y, with a product of 70 and a sum which is a maximum. @Mathematics

OpenStudy (anonymous):

I think the answer is -sqrt(70) for both numbers but I'm not really sure

OpenStudy (anonymous):

yes i think that has to be right, by sheer reason

OpenStudy (anonymous):

I took the derivative of 70/x + x then set it to 0

OpenStudy (anonymous):

you have to numbers and their product is 70 now you can write one as \[x\]and the other as \[\frac{70}{x}\] if you like and then try to maximize \[x+\frac{70}{x}\] if you like

OpenStudy (anonymous):

and that is what you will get. but you can also think. you made up the x and 70/x there is really no difference between them, by which i mean it is completely symmetric .

OpenStudy (anonymous):

that's true too

OpenStudy (anonymous):

so it has to have a max when they are equal

OpenStudy (anonymous):

i hope it is clear what the reasoning is, because calculus will compute the right answer for you but it is really more common sense

OpenStudy (anonymous):

well I took the min since they both need to be negative, does that make sense?

OpenStudy (anonymous):

actually it doesn't really matter in this case since there is only one critical number

OpenStudy (anonymous):

derivative is \[1-\frac{70}{x^2}\] and you can set this equal to zero and solve.

OpenStudy (anonymous):

but if you have to hand this in impress your teacher and say " the two numbers are x and y where x + y = 70 and since this is symmetric in x and y (switch x and y and you get the same equation) then the sum is max when they are equal

OpenStudy (anonymous):

I would have done that too! but this question is online for an online quiz Thank you very much for your help!